OpenStudy (anonymous):

Solve the inequality -3|6-x|less than -15 for x. Graph the solution.

OpenStudy (australopithecus):

-3|6-x|/-3 < -15/-3 = |6 - x|< 5 = -5 < |6-x| < 5 -5-6 <|-x| < 5 -6 -11/-1<-x/-1< -1/-1 flip the signs as you are dividing by a negative thus 11 > x > 1 thus you have that x is less than 11 but greater than 1 so the functions domain on the real line is as follows 

OpenStudy (australopithecus):

Note that thsi function goes infinitely close to 11 and 1 but never reaches those numbers

OpenStudy (anonymous):

thanks :)

OpenStudy (australopithecus):

to graph this remember you have two different functions, remember that the output of an abosolutve value always has to be positive. -3(-6+x) and -3(6 - x) Thus what ever is in the brackets has to be negative, a (-)(-) = + so check where what is inside the brackets is -6 + x = 0 x = 6 and we notice that if we go any higher than 6 the function will be give us a positive number and if we put anything in lower than 6 we get a negative (which we want). thus we have that for the first equation -3(-6+x), x < 6 we do the same for the second equation and we get -3(6 - x), x > 6 We notice also that if x = 6 the function is 0 because -3(0) = 0 so now you just sub the corresponding x values into the specific equation know the domain of the function is (1,11)

OpenStudy (australopithecus):

I hope this helps you with absolute values they are pretty intimidating at first, at least I felt that way when I first encountered them.